OpenStudy (anonymous):

you have 10 fewer quarters than dimes and 5 fewer nickels than quarters. the total value of the coins is 4.75. how many quarters, dimes and nickels do you have? please tell me how you set up the equations

OpenStudy (anonymous):

d-10=q q-5=n 25q+5n+10d=475

OpenStudy (anonymous):

when i do that, i get the equation 35d+5n=725..im stuck